OpenStudy (anonymous):

function of 2x+7y=-7 -4x-3y=-19

OpenStudy (kl0723):

solving for x?

OpenStudy (anonymous):

system of equations

OpenStudy (anonymous):

answer format is (a,b)

OpenStudy (kl0723):

in a system of equations just multiply the top part times 2: 4x +14y = -14 -4x + -3y = -19... cancel x values and get 11y = -33, y =-3 now plug in y value into equation #1 and solve for value of x... 2x + 7(-3) = -7... 2x = -7 + 21 2x = 14... x = 7 Answer = (7,-3)

OpenStudy (anonymous):

thank you!

OpenStudy (kl0723):

yw